# SegQuant: A Semantics-Aware and Generalizable Quantization Framework for Diffusion Models

### v3

@@ Introduction
To address these challenges, quantization (Gholami et al. 2021) has emerged as a practical solution for reducing the computational burden of diffusion inference. One promising approach to improving deployment efficiency is post-training quantization (Jacob et al. 2018) (PTQ), which reduces model precision without requiring retraining or extensive fine-tuning. Quantization significantly improves inference speed and memory usage. However, reducing numerical precision often leads to performance degradation, especially in complex models like diffusion models that rely on iterative refinement over multiple denoising steps. Therefore, developing effective quantization strategies that preserve generation fidelity while maximizing efficiency has become a critical research direction.
@@
Quantization has become a widely adopted technique for compressing deep models, with a majority of existing methods focusing on linear layers. These approaches can be broadly categorized into two types: weight-specific optimization and joint quantization of weights and activations. Weight-specific quantization techniques aim to minimize quantization error through mathematical optimization or blockwise reconstruction strategies (Nagel et al. 2020; Li et al. 2021; Frantar et al. 2022). In contrast, joint quantization methods consider both weights and activations during calibration, often leveraging smoothing or decomposition strategies to suppress outlier values (Xiao et al. 2023; Li* et al. 2025). Meanwhile, industrial quantization toolkits such as TensorRT (NVIDIA Corporation 2023) provide deployment-ready solutions, incorporating kernel-level optimizations and hardware-aware scheduling. While effective in many applications, these general-purpose toolkits often fall short when applied to inherently complex models like diffusion architectures, where domain-specific characteristics must be explicitly considered.
@@
The unique multi-step generative process and architectural complexity of diffusion models pose distinct challenges for quantization. Early efforts focused on timestep-dependent activation variance, proposing improved calibration schemes tailored to the denoising dynamics (Shang et al. 2023; Li et al. 2023). From an architectural perspective, several works have explored UNet-based diffusion models, designing quantization strategies that account for structural patterns such as concatenation, scaling, and temporal blocks (Li et al. 2023; Chu et al. 2024; Huang et al. 2024). More recently, attention has shifted to DiT-based diffusion models (Peebles and Xie 2023). For instance, PTQ4DiT (Wu et al. 2024) introduces channel-wise optimization through adaptive channel selection, while VIDIT-Q (Zhao et al. 2024) designs token-wise and timestep-wise techniques tailored to video generation. Despite these advances, most existing methods rely heavily on architecture-specific heuristics, limiting their generalizability across different diffusion variants and broader model families.
@@
where $\mathrm{round}(\cdot)$ denotes rounding the input to the nearest integer; $s$, $z$, and $[q_{\min},q_{\max}]$ represent the scale, zero-point, and quantization range, respectively. The original value $x$ can be approximately recovered by the inverse operation: $x\approx s\cdot(\hat{x}+z)$.
@@
Recent advancements have improved efficiency through techniques such as accumulating the denoising schedule (Song, Meng, and Ermon 2022) and leveraging vector fields (Lipman et al. 2023). Modern diffusion architectures often employ Transformer-based (Vaswani et al. 2017) backbones, exemplified by the DiT family of models (Peebles and Xie 2023), which offer superior scalability and performance compared to traditional UNet structures (Ronneberger, Fischer, and Brox 2015). DiT takes as input a latent feature map $\mathbf{z}$ obtained from a pretrained VAE (Kingma and Welling 2022), splits it into non-overlapping patches, and embeds them as tokens. It incorporates temporal information through learned positional embeddings and uses Adaptive LayerNorm (Perez et al. 2018) for conditional normalization. Cross-attention layers allow for flexible conditioning on text or other modalities. Finally, a linear layer predicts the noise $\boldsymbol{\epsilon}_{\theta}$, which is used to refine the sample in the next denoising step.
@@
In this section, we present SegQuant, a unified top-down quantization framework designed through an in-depth analysis of both diffusion-specific and general-purpose quantization techniques (Figure 1). In SegQuant, we organize existing quantization techniques into two modular components: the Optimizer (Xiao et al. 2023; Li* et al. 2025) and the Calibrator (Frantar et al. 2022), and further enhance it with our proposed modules, i.e., SegLinear and DualScale, which will be detailed in subsequent sections. In addition, the framework incorporates efficient CUDA kernels [^1] https://github.com/NVIDIA/cutlass for deployment, enabling acceleration without compromising the integrity of the quantization logic.
@@
To quantify this asymmetry, we analyzed activation outputs channel-wise across several representative layers from SD3.5-ControlNet using the COCO dataset (averaged over 30 timesteps). As shown in Table 1, a large proportion of channels consistently exhibit negative values. For instance, over 60%-70% of channels in certain AdaNorm and FFN layers remain predominantly negative during inference. However, conventional PTQ methods typically apply a single global scale (either symmetric or asymmetric) across the entire activation range. Due to the large spread of positive values, compression of the narrower negative range ([0.3,0][-0.3,0][ - 0.3 , 0 ]) is often too aggressive. Even with asymmetric quantization, which shifts the zero-point, the uniform distribution of bins leads to poor resolution for negative values. Figure 7 shows that modifying only the negative activations causes subtle but perceptible detail loss, particularly in texture and background consistency. This indicates that negative activations are not residual noise but semantically meaningful components. Although overall quality remains largely intact, this underscores preserving small negative values for high-fidelity generation.
@@
$=\left(s_{+}\cdot\mathbf{\hat{X}_{+}}+s_{-}\cdot\mathbf{\hat{X}_{-}}\right)\cdot\left(s_{w}\cdot\mathbf{\hat{W}}\right)$
$=s_{+}s_{w}\cdot\left(\mathbf{\hat{X}_{+}}\mathbf{\hat{W}}\right)+s_{-}s_{w}\cdot\left(\mathbf{\hat{X}_{-}}\mathbf{\hat{W}}\right),$
@@
We evaluate on COCO (Lin et al. 2015), and sample 5,000 images from MJHQ-30K (Li et al. 2024) and DCI (Urbanek et al. 2024), as suggested by recent work (Li* et al. 2025). Quantization quality is measured against FP16 outputs using FID (Heusel et al. 2017), LPIPS (Zhang et al. 2018), PSNR, SSIM (Wang et al. 2004), and Image Reward (Xu et al. 2023).
@@
We evaluate on three models: Stable Diffusion 3.5 Medium (Esser et al. 2024) (2B), FLUX.1-dev (Black Forest Labs 2024) (12B, DiT-based), and SDXL (Podell et al. 2023) (UNet-based). All experiments are run on NVIDIA RTX 4090 and L20 GPUs.
@@
SD3.5-DiT 8/8(int) Q-Diffusion 169.59169.59169.59 2.072-2.072- 2.072 0.6900.6900.690 7.927.927.92 0.2950.2950.295
PTQD 26.5326.5326.53 0.3090.3090.309 0.5200.5200.520 10.2010.2010.20 0.4170.4170.417
PTQ4DiT 16.4616.4616.46 0.7520.7520.752 0.4260.4260.426 12.1812.1812.18 0.5320.5320.532
TAC 17.7817.7817.78 0.7020.7020.702 0.4400.4400.440 11.9911.9911.99 0.5200.5200.520
SegQuant-A 13.9013.9013.90 0.9240.9240.924 0.3840.3840.384 12.7812.7812.78 0.5630.5630.563
SegQuant-G 12.3712.3712.37 0.8590.8590.859 0.3830.3830.383 12.8312.8312.83 0.5640.5640.564
4/8(int) PTQ4DiT 62.9862.9862.98 0.190-0.190- 0.190 0.5770.5770.577 10.0610.0610.06 0.4290.4290.429
SVDQuant 20.5820.5820.58 0.7250.7250.725 0.4560.4560.456 11.7611.7611.76 0.5230.5230.523
SegQuant-G 20.2220.2220.22 0.7620.7620.762 0.4530.4530.453 11.6911.6911.69 0.5210.5210.521
SDXL-UNet 8/8(int) Q-Diffusion 72.8172.8172.81 1.618-1.618- 1.618 0.5670.5670.567 11.9711.9711.97 0.3600.3600.360
PTQ4DiT 7.707.707.70 0.7360.7360.736 0.1910.1910.191 19.6619.6619.66 0.6910.6910.691
SegQuant-A 6.396.396.39 0.7750.7750.775 0.1410.1410.141 21.3321.3321.33 0.7420.7420.742
SegQuant-G 6.196.196.19 0.7640.7640.764 0.1340.1340.134 21.6021.6021.60 0.7500.7500.750
8/8(fp) Q-Diffusion 5.145.145.14 0.8970.8970.897 0.0930.0930.093 24.3124.3124.31 0.8270.8270.827
SegQuant-A 5.125.125.12 0.9010.9010.901 0.0930.0930.093 24.2824.2824.28 0.8270.8270.827
SegQuant-G 4.834.834.83 0.9030.9030.903 0.0820.0820.082 24.8424.8424.84 0.8380.8380.838
FLUX-DiT 8/8(int) Q-Diffusion 9.419.419.41 0.7320.7320.732 0.2990.2990.299 15.8715.8715.87 0.6330.6330.633
PTQ4DiT 11.9111.9111.91 0.6300.6300.630 0.3250.3250.325 15.3615.3615.36 0.6110.6110.611
SegQuant-A 5.895.895.89 0.8350.8350.835 0.1500.1500.150 19.8419.8419.84 0.7700.7700.770
SegQuant-G 5.565.565.56 0.8220.8220.822 0.1380.1380.138 20.3220.3220.32 0.7820.7820.782
4/8(int) PTQ4DiT 30.0930.0930.09 0.039-0.039- 0.039 0.5400.5400.540 12.0912.0912.09 0.5400.5400.540
SVDQuant 7.947.947.94 0.7830.7830.783 0.2320.2320.232 17.2917.2917.29 0.6970.6970.697
SegQuant-G 7.787.787.78 0.7890.7890.789 0.2250.2250.225 17.4817.4817.48 0.7020.7020.702
@@
SegQuant enhances quantization by integrating semantic-aware optimization and calibration. Optimization methods such as SmoothQuant (Xiao et al. 2023) and SVDQuant (Li* et al. 2025) utilize SegLinear with semantic segmentation for fine-grained tuning of the hyperparameter $\alpha$. For calibration, SegQuant adopts GPTQ (Frantar et al. 2022) or AMax, where SegLinear segments weights by output features to complement GPTQ. This design yields lower quantization error, as shown in Table 3. We further isolate the effects of SegLinear and DualScale, demonstrating their individual benefits to image quality in Table 4 and Fig. 9.
@@ Appendix B More Experimental Details and Results
SD3.5-DiT 8/8(int) Q-Diffusion 218.04218.04218.04 2.197-2.197- 2.197 0.3950.3950.395 16.0216.0216.02 0.7040.7040.704 8.218.218.21 0.3880.3880.388
PTQD 15.6215.6215.62 0.8220.8220.822 0.4340.4340.434 16.4716.4716.47 0.4540.4540.454 10.7910.7910.79 0.5230.5230.523
PTQ4DiT 14.2114.2114.21 0.9120.9120.912 0.4580.4580.458 16.3716.3716.37 0.4030.4030.403 12.6912.6912.69 0.5930.5930.593
TAC 16.8916.8916.89 0.8630.8630.863 0.4580.4580.458 16.3116.3116.31 0.4170.4170.417 12.5312.5312.53 0.5830.5830.583
SegQuant-A 10.9010.9010.90 1.0201.0201.020 0.4670.4670.467 16.3516.3516.35 0.3620.3620.362 13.2713.2713.27 0.6180.6180.618
SegQuant-G 11.0611.0611.06 0.9910.9910.991 0.4570.4570.457 16.3616.3616.36 0.3760.3760.376 13.1413.1413.14 0.6010.6010.601
4/8(int) PTQ4DiT 92.3692.3692.36 0.0850.0850.085 0.4110.4110.411 16.5116.5116.51 0.5700.5700.570 10.8310.8310.83 0.5040.5040.504
SVDQuant 25.4825.4825.48 0.8550.8550.855 0.4540.4540.454 16.4016.4016.40 0.4320.4320.432 12.3212.3212.32 0.5820.5820.582
SegQuant-G 26.7626.7626.76 0.8430.8430.843 0.4380.4380.438 16.4516.4516.45 0.4340.4340.434 12.4012.4012.40 0.5780.5780.578
SDXL-UNet 8/8(int) Q-Diffusion 74.8974.8974.89 1.574-1.574- 1.574 0.3960.3960.396 16.4816.4816.48 0.5370.5370.537 13.6813.6813.68 0.5170.5170.517
PTQ4DiT 7.517.517.51 0.6390.6390.639 0.4070.4070.407 16.7016.7016.70 0.2130.2130.213 19.1719.1719.17 0.7250.7250.725
SegQuant-A 5.885.885.88 0.6450.6450.645 0.4080.4080.408 16.7116.7116.71 0.1450.1450.145 21.3221.3221.32 0.7850.7850.785
SegQuant-G 5.725.725.72 0.6520.6520.652 0.4080.4080.408 16.7116.7116.71 0.1380.1380.138 21.6321.6321.63 0.7930.7930.793
8/8(fp) Q-Diffusion 4.944.944.94 0.8430.8430.843 0.4280.4280.428 16.4616.4616.46 0.1040.1040.104 23.3223.3223.32 0.8320.8320.832
SegQuant1 4.924.924.92 0.8440.8440.844 0.4280.4280.428 16.4616.4616.46 0.1040.1040.104 23.3123.3123.31 0.8320.8320.832
SegQuant2 4.634.634.63 0.8390.8390.839 0.4270.4270.427 16.4616.4616.46 0.0930.0930.093 23.9023.9023.90 0.8420.8420.842
FLUX-DiT 8/8(int) Q-Diffusion 8.378.378.37 0.8830.8830.883 0.4460.4460.446 16.5216.5216.52 0.3020.3020.302 15.2315.2315.23 0.6240.6240.624
PTQ4DiT 10.4710.4710.47 0.7160.7160.716 0.4600.4600.460 16.4516.4516.45 0.3280.3280.328 14.7314.7314.73 0.6200.6200.620
SegQuant-A 5.135.135.13 0.9070.9070.907 0.4450.4450.445 16.5216.5216.52 0.1550.1550.155 19.4019.4019.40 0.7700.7700.770
SegQuant-G 4.944.944.94 0.9000.9000.900 0.4440.4440.444 16.5416.5416.54 0.1430.1430.143 19.9419.9419.94 0.7840.7840.784
4/8(int) PTQ4DiT 34.9234.9234.92 0.0960.0960.096 0.4170.4170.417 16.6516.6516.65 0.5610.5610.561 11.6311.6311.63 0.5370.5370.537
SVDQuant 7.117.117.11 0.8800.8800.880 0.4360.4360.436 16.5716.5716.57 0.2420.2420.242 16.8516.8516.85 0.6930.6930.693
SegQuant-G 6.866.866.86 0.8820.8820.882 0.4390.4390.439 16.5616.5616.56 0.2320.2320.232 17.0317.0317.03 0.7000.7000.700
@@
[…]
SD3.5-DiT 8/8(int) Q-Diffusion 0.4440.4440.444 15.3115.3115.31 4/8(int) PTQ4DiT 0.428 15.86
PTQD 0.4410.4410.441 15.9115.9115.91
PTQ4DiT 0.4610.4610.461 15.9115.9115.91 SVDQuant 0.452 15.85
TAC 0.4600.4600.460 15.9015.9015.90
SegQuant-A 0.4680.4680.468 15.8615.8615.86
SegQuant-G 0.4660.4660.466 15.8815.8815.88 SegQuant-G 0.452 15.89
SDXL-UNet 8/8(int) Q-Diffusion 0.4000.4000.400 15.6115.6115.61 8/8(fp) Q-Diffusion 0.417 15.71
PTQ4DiT 0.4300.4300.430 15.7615.7615.76
SegQuant-A 0.4330.4330.433 15.7715.7715.77 SegQuant-A 0.4170.4170.417 15.7115.7115.71
SegQuant-G 0.4330.4330.433 15.7815.7815.78 SegQuant-G 0.4180.4180.418 15.7115.7115.71
FLUX-DiT 8/8(int) Q-Diffusion 0.4440.4440.444 15.9815.9815.98 4/8(int) PTQ4DiT 0.416 15.98
PTQ4DiT 0.4610.4610.461 15.8815.8815.88
SegQuant-A 0.4400.4400.440 15.9415.9415.94 SVDQuant 0.4350.4350.435 15.9915.9915.99
SegQuant-G 0.4400.4400.440 15.9315.9315.93 SegQuant-G 0.4370.4370.437 16.0116.0116.01

### v4

@@ Introduction
To address these challenges, quantization (Gholami et al. 2022) has emerged as a practical solution for reducing the computational burden of diffusion inference. One promising approach to improving deployment efficiency is post-training quantization (Jacob et al. 2018) (PTQ), which reduces model precision without requiring retraining or extensive fine-tuning. Quantization significantly improves inference speed and memory usage. However, reducing numerical precision often leads to performance degradation, especially in complex models like diffusion models that rely on iterative refinement over multiple denoising steps. Therefore, developing effective quantization strategies that preserve generation fidelity while maximizing efficiency has become a critical research direction.
@@
Quantization has become a widely adopted technique for compressing deep models, with a majority of existing methods focusing on linear layers. These approaches can be broadly categorized into two types: weight-specific optimization and joint quantization of weights and activations. Weight-specific quantization techniques aim to minimize quantization error through mathematical optimization or blockwise reconstruction strategies (Nagel et al. 2020; Li et al. 2021; Frantar et al. 2023). In contrast, joint quantization methods consider both weights and activations during calibration, often leveraging smoothing or decomposition strategies to suppress outlier values (Xiao et al. 2023; Li* et al. 2025). Meanwhile, industrial quantization toolkits such as TensorRT (NVIDIA Corporation 2023) provide deployment-ready solutions, incorporating kernel-level optimizations and hardware-aware scheduling. While effective in many applications, these general-purpose toolkits often fall short when applied to inherently complex models like diffusion architectures, where domain-specific characteristics must be explicitly considered.
@@
The unique multi-step generative process and architectural complexity of diffusion models pose distinct challenges for quantization. Early efforts focused on timestep-dependent activation variance, proposing improved calibration schemes tailored to the denoising dynamics (Shang et al. 2023; Li et al. 2023). From an architectural perspective, several works have explored UNet-based diffusion models, designing quantization strategies that account for structural patterns such as concatenation, scaling, and temporal blocks (Li et al. 2023; Chu et al. 2024; Huang et al. 2024). More recently, attention has shifted to DiT-based diffusion models (Peebles and Xie 2023). For instance, PTQ4DiT (Wu et al. 2024) introduces channel-wise optimization through adaptive channel selection, while VIDIT-Q (Zhao et al. 2025) designs token-wise and timestep-wise techniques tailored to video generation. Despite these advances, most existing methods rely heavily on architecture-specific heuristics, limiting their generalizability across different diffusion variants and broader model families.
@@
where $\mathrm{round}(\cdot)$ denotes rounding the input to the nearest integer; $s$, $z$, and $q_{\min},q_{\max}]$ represent the scale, zero-point, and quantization range, respectively. The original value $x$ can be approximately recovered by the inverse operation: $x\approx s\cdot(\hat{x}+z)$.
@@
Recent advancements have improved efficiency through techniques such as accumulating the denoising schedule (Song, Meng, and Ermon 2021) and leveraging vector fields (Lipman et al. 2023). Modern diffusion architectures often employ Transformer-based (Vaswani et al. 2017) backbones, exemplified by the DiT family of models (Peebles and Xie 2023), which offer superior scalability and performance compared to traditional UNet structures (Ronneberger, Fischer, and Brox 2015). DiT takes as input a latent feature map $\mathbf{z}$ obtained from a pretrained VAE (Kingma and Welling 2022), splits it into non-overlapping patches, and embeds them as tokens. It incorporates temporal information through learned positional embeddings and uses Adaptive LayerNorm (Perez et al. 2018) for conditional normalization. Cross-attention layers allow for flexible conditioning on text or other modalities. Finally, a linear layer predicts the noise $\boldsymbol{\epsilon}_{\theta}$, which is used to refine the sample in the next denoising step.
@@
In this section, we present SegQuant, a unified top-down quantization framework designed through an in-depth analysis of both diffusion-specific and general-purpose quantization techniques (Figure 1). In SegQuant, we organize existing quantization techniques into two modular components: the Optimizer (Xiao et al. 2023; Li* et al. 2025) and the Calibrator (Frantar et al. 2023), and further enhance it with our proposed modules, i.e., SegLinear and DualScale, which will be detailed in subsequent sections. In addition, the framework incorporates efficient CUDA kernels [^1] https://github.com/NVIDIA/cutlass for deployment, enabling acceleration without compromising the integrity of the quantization logic.
@@
To quantify this asymmetry, we analyzed activation outputs channel-wise across several representative layers from SD3.5-ControlNet using the COCO dataset (averaged over 30 timesteps). As shown in Table 1, a large proportion of channels consistently exhibit negative values. For instance, over 60%-70% of channels in certain AdaNorm and FFN layers remain predominantly negative during inference. However, conventional PTQ methods typically apply a single global scale (either symmetric or asymmetric) across the entire activation range. Due to the large spread of positive values, compression of the narrower negative range ([0.3,0][-0.3,0]) is often too aggressive. Even with asymmetric quantization, which shifts the zero-point, the uniform distribution of bins leads to poor resolution for negative values. Figure 7 shows that modifying only the negative activations causes subtle but perceptible detail loss, particularly in texture and background consistency. This indicates that negative activations are not residual noise but semantically meaningful components. Although overall quality remains largely intact, this underscores preserving small negative values for high-fidelity generation.
@@
[…]
@@
We evaluate on COCO (Lin et al. 2014), and sample 5,000 images from MJHQ-30K (Li et al. 2024) and DCI (Urbanek et al. 2024), as suggested by recent work (Li* et al. 2025). Quantization quality is measured against FP16 outputs using FID (Heusel et al. 2017), LPIPS (Zhang et al. 2018), PSNR, SSIM (Wang et al. 2004), and Image Reward (Xu et al. 2023).
@@
We evaluate on three models: Stable Diffusion 3.5 Medium (Esser et al. 2024) (2B), FLUX.1-dev (Black Forest Labs 2024) (12B, DiT-based), and SDXL (Podell et al. 2024) (UNet-based). All experiments are run on NVIDIA RTX 4090 and L20 GPUs.
@@
SD3.5-DiT 8/8(int) Q-Diffusion $169.59$ 2.072-2.072 $0.690$ $7.92$ $0.295$
PTQD $26.53$ $0.309$ $0.520$ $10.20$ $0.417$
PTQ4DiT $16.46$ $0.752$ $0.426$ $12.18$ $0.532$
TAC $17.78$ $0.702$ $0.440$ $11.99$ $0.520$
SegQuant-A $13.90$ $0.924$ $0.384$ $12.78$ $0.563$
SegQuant-G $12.37$ $0.859$ $0.383$ $12.83$ $0.564$
4/8(int) PTQ4DiT $62.98$ 0.190-0.190 $0.577$ $10.06$ $0.429$
SVDQuant $20.58$ $0.725$ $0.456$ $11.76$ $0.523$
SegQuant-G $20.22$ $0.762$ $0.453$ $11.69$ $0.521$
SDXL-UNet 8/8(int) Q-Diffusion $72.81$ 1.618-1.618 $0.567$ $11.97$ $0.360$
PTQ4DiT $7.70$ $0.736$ $0.191$ $19.66$ $0.691$
SegQuant-A $6.39$ $0.775$ $0.141$ $21.33$ $0.742$
SegQuant-G $6.19$ $0.764$ $0.134$ $21.60$ $0.750$
8/8(fp) Q-Diffusion $5.14$ $0.897$ $0.093$ $24.31$ $0.827$
SegQuant-A $5.12$ $0.901$ $0.093$ $24.28$ $0.827$
SegQuant-G $4.83$ $0.903$ $0.082$ $24.84$ $0.838$
FLUX-DiT 8/8(int) Q-Diffusion $9.41$ $0.732$ $0.299$ $15.87$ $0.633$
PTQ4DiT $11.91$ $0.630$ $0.325$ $15.36$ $0.611$
SegQuant-A $5.89$ $0.835$ $0.150$ $19.84$ $0.770$
SegQuant-G $5.56$ $0.822$ $0.138$ $20.32$ $0.782$
4/8(int) PTQ4DiT $30.09$ 0.039-0.039 $0.540$ $12.09$ $0.540$
SVDQuant $7.94$ $0.783$ $0.232$ $17.29$ $0.697$
SegQuant-G $7.78$ $0.789$ $0.225$ $17.48$ $0.702$
@@
SegQuant enhances quantization by integrating semantic-aware optimization and calibration. Optimization methods such as SmoothQuant (Xiao et al. 2023) and SVDQuant (Li* et al. 2025) utilize SegLinear with semantic segmentation for fine-grained tuning of the hyperparameter $\alpha$. For calibration, SegQuant adopts GPTQ (Frantar et al. 2023) or AMax, where SegLinear segments weights by output features to complement GPTQ. This design yields lower quantization error, as shown in Table 3. We further isolate the effects of SegLinear and DualScale, demonstrating their individual benefits to image quality in Table 4 and Fig. 9.
@@ Appendix B More Experimental Details and Results
SD3.5-DiT 8/8(int) Q-Diffusion $218.04$ 2.197-2.197 $0.395$ $16.02$ $0.704$ $8.21$ $0.388$
PTQD $15.62$ $0.822$ $0.434$ $16.47$ $0.454$ $10.79$ $0.523$
PTQ4DiT $14.21$ $0.912$ $0.458$ $16.37$ $0.403$ $12.69$ $0.593$
TAC $16.89$ $0.863$ $0.458$ $16.31$ $0.417$ $12.53$ $0.583$
SegQuant-A $10.90$ $1.020$ $0.467$ $16.35$ $0.362$ $13.27$ $0.618$
SegQuant-G $11.06$ $0.991$ $0.457$ $16.36$ $0.376$ $13.14$ $0.601$
4/8(int) PTQ4DiT $92.36$ $0.085$ $0.411$ $16.51$ $0.570$ $10.83$ $0.504$
SVDQuant $25.48$ $0.855$ $0.454$ $16.40$ $0.432$ $12.32$ $0.582$
SegQuant-G $26.76$ $0.843$ $0.438$ $16.45$ $0.434$ $12.40$ $0.578$
SDXL-UNet 8/8(int) Q-Diffusion $74.89$ 1.574-1.574 $0.396$ $16.48$ $0.537$ $13.68$ $0.517$
PTQ4DiT $7.51$ $0.639$ $0.407$ $16.70$ $0.213$ $19.17$ $0.725$
SegQuant-A $5.88$ $0.645$ $0.408$ $16.71$ $0.145$ $21.32$ $0.785$
SegQuant-G $5.72$ $0.652$ $0.408$ $16.71$ $0.138$ $21.63$ $0.793$
8/8(fp) Q-Diffusion $4.94$ $0.843$ $0.428$ $16.46$ $0.104$ $23.32$ $0.832$
SegQuant1 $4.92$ $0.844$ $0.428$ $16.46$ $0.104$ $23.31$ $0.832$
SegQuant2 $4.63$ $0.839$ $0.427$ $16.46$ $0.093$ $23.90$ $0.842$
FLUX-DiT 8/8(int) Q-Diffusion $8.37$ $0.883$ $0.446$ $16.52$ $0.302$ $15.23$ $0.624$
PTQ4DiT $10.47$ $0.716$ $0.460$ $16.45$ $0.328$ $14.73$ $0.620$
SegQuant-A $5.13$ $0.907$ $0.445$ $16.52$ $0.155$ $19.40$ $0.770$
SegQuant-G $4.94$ $0.900$ $0.444$ $16.54$ $0.143$ $19.94$ $0.784$
4/8(int) PTQ4DiT $34.92$ $0.096$ $0.417$ $16.65$ $0.561$ $11.63$ $0.537$
SVDQuant $7.11$ $0.880$ $0.436$ $16.57$ $0.242$ $16.85$ $0.693$
SegQuant-G $6.86$ $0.882$ $0.439$ $16.56$ $0.232$ $17.03$ $0.700$
@@
$uparrow C.IQA\uparrow C.SCR\uparrow$
SD3.5-DiT 8/8(int) Q-Diffusion $0.444$ $15.31$ 4/8(int) PTQ4DiT 0.428 15.86
PTQD $0.441$ $15.91$
PTQ4DiT $0.461$ $15.91$ SVDQuant 0.452 15.85
TAC $0.460$ $15.90$
SegQuant-A $0.468$ $15.86$
SegQuant-G $0.466$ $15.88$ SegQuant-G 0.452 15.89
SDXL-UNet 8/8(int) Q-Diffusion $0.400$ $15.61$ 8/8(fp) Q-Diffusion 0.417 15.71
PTQ4DiT $0.430$ $15.76$
SegQuant-A $0.433$ $15.77$ SegQuant-A $0.417$ $15.71$
SegQuant-G $0.433$ $15.78$ SegQuant-G $0.418$ $15.71$
FLUX-DiT 8/8(int) Q-Diffusion $0.444$ $15.98$ 4/8(int) PTQ4DiT 0.416 15.98
PTQ4DiT $0.461$ $15.88$
SegQuant-A $0.440$ $15.94$ SVDQuant $0.435$ $15.99$
SegQuant-G $0.440$ $15.93$ SegQuant-G $0.437$ $16.01$